How many grams of a
14.8% sugar solution contain 55.0 g of sugar?
1
Expert's answer
2019-10-02T08:02:30-0400

Mass concentration can be expressed as:


"w=\frac{m_{sugar}}{m_{solution}}*100\%"


Thus, the mass of solution is:


"m_{solution}=\frac{m_{sugar}*100\%}{w}=\frac{55.0g*100\%}{14.8\%}=371.6g"


Finally, we can say that 371.6 g of a given solution contains 55.0 g of sugar.